solve x^2=y+10
3x-y=9
1
Expert's answer
2012-04-17T09:56:00-0400
from second equation y=3x-9
from first
x^2=y+10=3x-9+10=3x+1
x^2-3x+1=0;
x1,2=(3+-sqrt(5))/2;
y1,2=3x(1,2)-9